Find an equation where ax^2+bx+c=0 with solutions of 2/5 and -5.
Answer by nerdybill(7384)

since you have solutions of 2/5 and -5
two factors must be:
(x - 2/5) and (x+5)
.
FOIL the two factors to get the polynomial:
(x - 2/5)(x+5) = 0
x^2 + 5x -2x/5 - 2 = 0
x^2 + 25x/5 -2x/5 - 2 = 0
x^2 + 23x/5 - 2 = 0
x^2 + (23/5)x - 2 = 0